# AlphaDou: High-Performance End-to-End Doudizhu AI Integrating Bidding

### v1

@@ Abstract
Artificial intelligence for card games has long been a popular topic in AI research. In recent years, complex card games like Mahjong and Texas Hold’em have been solved, with corresponding AI programs reaching the level of human experts. However, the game of Dou Di Zhu presents significant challenges due to its vast state/action space and unique characteristics involving reasoning about competition and cooperation, making the game extremely difficult to solve.The RL model DouZero, trained using the Deep Monte Carlo algorithm framework, has shown excellent performance in DouDiZhu. However, there are differences between its simplified game environment and the actual Dou Di Zhu environment, and its performance is still a considerable distance from that of human experts. This paper modifies the Deep Monte Carlo algorithm framework by using reinforcement learning to obtain a neural network that simultaneously estimates win rates and expectations. The action space is pruned using expectations, and strategies are generated based on win rates. This RL model is trained in a realistic DouDiZhu environment and achieves a state-of-the-art level among publicly available models.
Introduction
Games can be broadly classified into two categories: perfect-information games (PIGs) and imperfect-information games (IIGs). In PIGs, players can observe all game states, such as in Shogi, Go, and Chess. In contrast, IIGs involve scenarios where participants cannot access complete information about other players, such as in heads-up Texas Hold’em. Reinforcement learning (RL) has been successfully applied to create numerous game AIs. RL algorithms have achieved remarkable success in both PIGs and IIGs, exemplified by AlphaGo (Silver et al. 2016) and AlphaZero (Silver et al. 2017) in Go, AlphaStar (Vinyals et al. 2019) in StarCraft II, OpenAI Five (OpenAI et al. 2019) in Dota 2, Suphx (Li et al. 2020) in Mahjong, Douzero (Zha et al. 2021) in Doudizhu, NukkiAI (Bouzy, Rimbaud, and Ventos 2020) in Contract Bridge, and AlphaHoldem (Zhao et al. 2022a) in Hold’em.
@@ Introduction
This work aims to develop a high-performance end-to-end Doudizhu AI model that incorporates bidding. Doudizhu, also known as Fighting the Landlord, is the most popular card game in China. Doudizhu is a 3-player IIG where players bid based on their hands, and the winning bidder becomes the Landlord. The remaining players form the Peasants team to oppose the Landlord. If any Peasant player wins, the entire team wins. The Landlord wins double rewards, while each Peasant player receives a single reward if the team wins, and vice versa. Rewards are related to the bid score and the occurrence of "bombs" (four cards of the same rank) or "rockets" during the game. Players with good hands tend to bid high to become the Landlord for higher returns. Moreover, Doudizhu has a large, flexible, and diverse action space with thousands of possible states ($10^{83}$) and actions (27,472) due to card combinations and complex rules (Zha et al. 2019). Additionally, rewards in Doudizhu are sparse and highly variable, only awarded at the end of the game and influenced by the bidding phase, the number of "bombs" during the game, and "spring" rewards. These characteristics make training a Doudizhu AI extremely challenging, and existing Doudizhu AIs exhibit certain issues.
@@
Previous research on Doudizhu AI has primarily focused on the playing phase, neglecting the bidding phase or employing completely random bid strategies. DeltaDou (Jiang et al. 2019) is the first AI program to achieve human-level performance compared to top human players, using an AlphaZero-like algorithm with Bayesian methods to infer hidden information and sample other players’ actions based on their policy networks. However, the vast action space in Doudizhu limits DeltaDou’s effectiveness. Douzero introduced Deep Monte Carlo (DMC), which combines the conventional Monte Carlo method with deep neural networks. In a Monte Carlo self-play framework, deep neural networks first estimate the value of each action (Q-value), then select the action with the highest Q-value as a training label or final move. DMC addresses the challenge of Doudizhu’s large action space, making training more stable. Doudizhu with DMC successfully outperformed other RL algorithms, including Deep-Q-Learning (DQN) (Mnih et al. 2015; You et al. 2019), Combination Q-Network (CQN) (You et al. 2019), and A3C (Mnih et al. 2016; You et al. 2019). Subsequently, (Wang, Wu, and Lai 2023) noted that the score distribution in gambling games is a combination of winning and losing score distributions, with risk-averse strategies resulting in a significant gap between them. Previous value-based methods directly predicted this combined distribution, leading to high variance and unstable training. They proposed WagerWin (Wang, Wu, and Lai 2023), which introduces probability and value factorization, enabling individual updates of the winning probability, losing Q-value, and winning Q-value. This method stabilizes the training of gambling game AIs. However, WagerWin primarily accelerated AI convergence without significantly improving the optimal policy and winning rate. In addition, several variants of DouZero have been proposed, such as DouZero+ (Zhao et al. 2022b) and Full DouZero+ (Zhao et al. 2024), which incorporate predictions of the opponents’ hands based on the original DouZero model. However, these variants do not provide quantifiable improvements over the original DouZero. Mdou (Luo et al. 2024), on the other hand, consolidates the three models of DouZero into a single model for training, resulting in faster convergence. Despite this, the overall performance of the model does not show significant improvement compared to DouZero. RARSMSDou (Luo and Tan 2024) utilized the PPO framework (Schulman et al. 2017) to enhance Doudizhu AI, addressing the large action space by abstracting actions into several major categories, training a PPO model to select categories, and then training a DMC model to choose actions within the selected category. RARSMSDou outperformed DouZero.
In this paper, we propose AlphaDou, an end-to-end Doudizhu AI system that incorporates bidding. Our model does not require abstract state/action spaces or any human knowledge. By utilizing more information, our model can adjust its playing strategy based on bidding results. We conducted several experiments, and our results indicate that our bid strategy surpasses the bid network using supervised learning with human expert data. Also, Our model outperforms existing Doudizhu AIs regardless of whether the bid strategy is applied.
The Game of Doudizhu
@@
Cardplay
During the Cardplay phase, players take turns playing cards. Each game consists of multiple rounds, starting with a player playing a valid card combination (e.g., solo, pair). The first round is initiated by the Landlord. Subsequent players must either pass or defeat the previous hand by playing a higher-ranked combination (an action has a rank, refer to Appendix AA). The round continues until two consecutive players pass. Then, the player who played the last hand starts the next round. The objective is to clear all cards from one’s hand to win. Each "Bomb" and "Rocket" can double the game’s stakes. If the Landlord wins, they receive double the rewards, whereas if the Peasant team wins, each Peasant player receives single rewards. Rewards are influenced by the bid score and the presence of Bombs or Rockets. Bombs surpass any action. The only way to defeat a Bomb is with a higher-ranked Bomb or a Rocket. The Rocket is the highest action in the game and can beat any Bomb or action. When a Bomb or Rocket is played, the points at stake double. For instance, if the winning bid is 3 points at the start, it becomes 6 points if a Bomb is played and 12 points if another Bomb is played. With two Bombs played, the Landlord stands to win/lose 24 points, and each Peasant stands to win/lose 12 points. The game concludes when a player clears all their cards. To encourage more aggressive play, Doudizhu includes a "Spring" reward: if throughout the game, the Peasant team makes no plays other than passing, or the Landlord only passes once, it is termed as Spring or Anti-Spring, respectively, doubling the reward (equivalent to playing an additional Bomb).
@@
Card Representation and Neural Architecture
@@
Deep Monte-Carlo
Monte Carlo (MC) methods are a class of methods that estimate strategies and value functions by modeling sample paths. Monte Carlo methods are very effective in episodic tasks to estimate the value function by taking every-visit MC approach (Sutton 1998).
@@
DMC with Probability and Value Factorization
Considering the significant gap between the distribution of winning scores and losing scores, we perform Value Factorization on the Q-value (Wang, Wu, and Lai 2023). Given that there are no ties in Doudizhu and the outcomes of the game (win or lose) are mutually exclusive, we have:
@@
We use the epsilon-greedy method to introduce exploration into the strategy $\pi(s)$ used to generate data. In appendix BB, we utilized the win rate model douzero-wp and the expected value model douzero-adp provided by DouZero to verify that our proposed strategy generation method outperforms the "choosing the move with the highest expectation" strategy.
EXPERIMENTS
In this chapter, we compare the performance of the AlphaDou card-playing model (CardModel) with Douzero and Douzero Resnet. Douzero Resnet is the state-of-the-art Dou Dizhu AI based on the Douzero algorithm, replacing the LSTM neural network in Douzero with ResNet, significantly improving performance compared to Douzero. The weights and code for Douzero Resnet are open-sourced at https://github.com/Vincentzyx/Douzero_Resnet. We also compare the AlphaDou bid model (Bid Model) with a supervised learning bid model (Douzero Resnet Bid). The Douzero Resnet Bid we used is derived from Douzero Resnet: fixing the landlord player’s hand, randomly distributing 1000 sets of farmer hands and landlord hands, loading the Douzero Resnet model for games, obtaining a mean score from the results of 1000 games, using the hand as input, and the mean score as the label for supervised learning. When applying the model, a threshold is set for bidding: a model output greater than -0.1 bids 1 point, greater than 0 bids 2 points, and greater than 0.1 bids 3 points. The Douzero demonstration website https://www.douzero.org/bid also has a bidding model, but its bidding method is not based on a 3-point system, and it does not have an interface reserved for testing. Our AI system is trained on a server with 4 Intel(R) Xeon(R) Gold 6330 CPUs @ 2.10GHz and a GeForce RTX 4090 GPU in the Ubuntu 20.04 operating system. The code is available in the supplement.
Compare Bid Model to Douzero Resnet Bid
@@
Metrics. Following (Jiang et al. 2019), given an algorithm A and an opponent B, we use two metrics to compare the performance of A and B:
@@
Compare Card Model to Benchmarks with Random Bidding Phase
To evaluate the performance of the Card Model, we followed the approach of (Jiang et al. 2019) and douzero (Zha et al. 2021), initiating a competition between the Landlord and the Peasants. We reduce variance by playing each deck twice. Specifically, for two competing algorithms A and B, they will first play with A as the Landlord and B as the Peasants for a given deck. Then, they swap roles, with A as the Peasants and B as the Landlord, and play the same deck again. A total of 4,000 games were conducted. Considering that the Bid result is random, we set the initial score of the game to 2 points for the Landlord’s win and 1 point for the Peasants’ win, with each bomb doubling the final score (we define this scoring method as ADP1). The Card Model needs to decide the playing strategy based on the bidding process, and the random bidding process will lead to a decline in model performance because the random testing deck distribution deviates from the training process.
@@ A B Card Model Douzero Resnet Douzero
Card Model - - 0.5224 0.1032 0.5970 0.4343
Douzero Resnet -0.4776 -0.1032 - - 0.5702 0.2688
Douzero 0.4030 -0.4343 0.4298 -0.2688 - -
@@
Compare Card Model to benchmarks with Bidding Phase
@@
Card Model - - - 0.5442 0.3150 0.7376 0.6167 0.5759 1.5852
Douzero Resnet 0.4558 -0.3150 -0.7376 - - - 0.5809 0.3140 0.9371
Douzero 0.3833 -0.5759 -1.5852 0.4191 -0.3140 -0.9371 - - -
@@
Card Model 0.5145 -0.0483 0.7850 4.6451
Douzero Resnet 0.4903 -0.2088 0.7709 4.2965
@@
Case Study: Bid Model vs Douzero Resnet Bid
Compared to threshold bidding, reinforcement learning bidding has a more flexible handling of different bidding situations. Here, we analyze a hand with the cards 333444569TTJJQKK2. The hand score given by Douzero Resnet Bid is -0.987, indicating that this hand is very weak. Firstly, the only high card is 2, and the absence of 7, 8, A, B, and R means that there is a high probability that other players’ hands form bombs. Using Douzero Resnet Bid, the choice would be “0 points”. For the same hand, the Bid Model gives different bidding scores based on the bidding order. In different situations, the Bid Model’s bidding strategy varies, as shown in Table 8. The Bid Model tends to bid 2 points in the first position, 0 points in the second position, and 3 points in the third position (only if 2 points were bid in the first position).
In the first position, the model bids 2 points, but the predicted winning and losing points differ significantly from the results in the last column of the table. In the table, the Landlord wins with 2 points. However, when bidding 2 points in the first position, the model tends to predict that the Landlord will eventually be outbid by other players with 3 points. Although the hand is weak, the presence of the airplane 333444 makes the card type look tidy. This implies that other players will have fewer small cards. When a player has very few small cards and many high cards, bidding 3 points can yield the greatest value. Even if the first position bids 2 points, they won’t become the final Landlord. If other players don’t feel confident enough to bid 3 points, it means that the high cards are either with the Landlord or with other players but not in a neat formation, requiring many turns to play out. In this case, a 2-point Landlord can still win, preventing a draw and retaining the potential to exploit weaker hands with weak card power.
In the third position, the model tends to bid 3 points, with $Q_{w}$ approaching 9 points. The model believes that high card power (Rocket) is in the Landlord’s hand, and the remaining players’ card types are not neat, making it difficult to handle the airplane card type (333444) and potential consecutive pairs (99TTJJ, TTJJQQKK, etc.). In this case, bidding 3 points to become the Landlord could result in gaining a bomb or rocket along with a Spring, cleverly utilizing the bidding position to allow weak card power to exploit even weaker card power.
Position First Second Agent Win rate $Q_{w}$ $Q_{l}$ $Q$
First / / 2 0.2724 4.4712 -5.7912 -2.9976
Second 0 / 0 0.1122 3.7488 -5.8030 -4.7328
Second 1 / 0 0.1203 3.8616 -6.0600 -4.8672
Second 2 / 0 0.2407 3.3096 -4.3200 -2.4816
Third 0 0 3 0.5154 9.3168 -7.7808 1.0296
Third 0 1 3 0.4742 9.4392 -7.8432 0.3528
Third 0 2 3 0.4378 8.5080 -7.5552 -0.5232
Third 1 0 3 0.4907 9.4536 -7.6920 0.7200
Third 1 2 3 0.4281 7.9008 -7.1016 -0.6792
Third 2 0 0 0.2477 2.2656 -2.7432 -1.5024
Case Study: Card Model vs Douzero Resnet vs Douzero
Figure 3: Play scenarios after the landlord regains card rights
Figure 4: Second critical point in the game
Conclusion

### v2

@@ Abstract
Artificial intelligence for card games has long been a popular topic in AI research. In recent years, complex card games like Mahjong and Texas Hold’em have been solved, with corresponding AI programs reaching the level of human experts. However, the game of Doudizhu presents significant challenges due to its vast state/action space and unique characteristics involving reasoning about competition and cooperation, making the game extremely difficult to solve.The RL model Douzero, trained using the Deep Monte Carlo algorithm framework, has shown excellent performance in Doudizhu. However, there are differences between its simplified game environment and the actual Doudizhu environment, and its performance is still a considerable distance from that of human experts. This paper modifies the Deep Monte Carlo algorithm framework by using reinforcement learning to obtain a neural network that simultaneously estimates win rates and expectations. The action space is pruned using expectations, and strategies are generated based on win rates. The modified algorithm enables the AI to perform the full range of tasks in the Doudizhu game, including bidding and cardplay. The model was trained in a actual Doudizhu environment and achieved state-of-the-art performance among publicly available models. We hope that this new framework will provide valuable insights for AI development in other bidding-based games.
1 Introduction
Games can be broadly classified into two categories: perfect-information games (PIGs) and imperfect-information games (IIGs). In PIGs, players can observe all game states, such as in Shogi, Go, and Chess. In contrast, IIGs involve scenarios where participants cannot access complete information about other players, such as in heads-up Texas Hold’em. Reinforcement learning (RL) has been successfully applied to create numerous game AIs. RL algorithms have achieved remarkable success in both PIGs and IIGs, exemplified by AlphaGo (Silver et al., 2016) and AlphaZero (Silver et al., 2017) in Go, AlphaStar (Vinyals et al., 2019) in StarCraft II, OpenAI Five (OpenAI et al., 2019) in Dota 2, Suphx (Li et al., 2020) in Mahjong, Douzero (Zha et al., 2021) in Doudizhu, NukkiAI (Bouzy et al., 2020) in Contract Bridge, and AlphaHoldem (Zhao et al., 2022a) in Hold’em.
@@ 1 Introduction
This work aims to develop a high-performance end-to-end Doudizhu AI model that incorporates bidding. Doudizhu, also known as Fighting the Landlord, is the most popular card game in China. Doudizhu is a 3-player IIG where players bid based on their hands, and the winning bidder becomes the Landlord. The remaining players form the Peasants team to oppose the Landlord. If any Peasant player wins, the entire team wins. The Landlord wins double rewards, while each Peasant player receives a single reward if the team wins, and vice versa. Rewards are related to the bid score and the occurrence of "bombs" (four cards of the same rank) or "rockets" during the game. Players with good hands tend to bid high to become the Landlord for higher returns. Moreover, Doudizhu has a large, flexible, and diverse action space with thousands of possible states ($10^{83}$) and actions (27,472) due to card combinations and complex rules (Zha et al., 2019). Additionally, rewards in Doudizhu are sparse and highly variable, only awarded at the end of the game and influenced by the bidding phase, the number of "bombs" during the game, and "spring" rewards. These characteristics make training a Doudizhu AI extremely challenging, and existing Doudizhu AIs exhibit certain issues.
@@
Previous research on Doudizhu AI has primarily focused on the playing phase, neglecting the bidding phase or employing completely random bid strategies. DeltaDou (Jiang et al., 2019) is the first AI program to achieve human-level performance compared to top human players, using an AlphaZero-like algorithm with Bayesian methods to infer hidden information and sample other players’ actions based on their policy networks. However, the vast action space in Doudizhu limits DeltaDou’s effectiveness. Douzero introduced Deep Monte Carlo (DMC), which combines the conventional Monte Carlo method with deep neural networks. In a Monte Carlo self-play framework, deep neural networks first estimate the value of each action (Q-value), then select the action with the highest Q-value as a training label or final move. DMC addresses the challenge of Doudizhu’s large action space, making training more stable. Doudizhu with DMC successfully outperformed other RL algorithms, including Deep-Q-Learning (DQN) (Mnih et al., 2015; You et al., 2019), Combination Q-Network (CQN) (You et al., 2019), and A3C (Mnih et al., 2016; You et al., 2019). Subsequently, (Wang et al., 2023) noted that the score distribution in gambling games is a combination of winning and losing score distributions, with risk-averse strategies resulting in a significant gap between them. Previous value-based methods directly predicted this combined distribution, leading to high variance and unstable training. They proposed WagerWin (Wang et al., 2023), which introduces probability and value factorization, enabling individual updates of the winning probability, losing Q-value, and winning Q-value. This method stabilizes the training of gambling game AIs. However, WagerWin primarily accelerated AI convergence without significantly improving the optimal policy and winning rate. In addition, several variants of Douzero have been proposed, such as Douzero+ (Zhao et al., 2022b) and Full Douzero+ (Zhao et al., 2024), which incorporate predictions of the opponents’ hands based on the original Douzero model. However, these variants do not provide quantifiable improvements over the original Douzero. Mdou (Luo et al., 2024), on the other hand, consolidates the three models of Douzero into a single model for training, resulting in faster convergence. Despite this, the overall performance of the model does not show significant improvement compared to Douzero. RARSMSDou (Luo & Tan, 2024) utilized the PPO framework (Schulman et al., 2017) to enhance Doudizhu AI, addressing the large action space by abstracting actions into several major categories, training a PPO model to select categories, and then training a DMC model to choose actions within the selected category. RARSMSDou outperformed Douzero.
In this paper, we introduce AlphaDou, an end-to-end DouDiZhu AI system that integrates bidding. Our model eliminates the need for abstract state/action spaces or any human-crafted knowledge. It simultaneously estimates both the win rate and the expected value of a given state, enabling it to prune alternative moves based on expectations and select the optimal move strategy based on win rates. Moreover, the model is capable of perceiving bidding outcomes and dynamically adjusting its move strategy accordingly. Extensive experiments demonstrate that our bidding strategy surpasses the performance of bidding networks trained via supervised learning. Additionally, during the cardplay phase, our model consistently outperforms existing DouDiZhu AI systems, whether or not a bidding strategy is employed. The training code for AlphaDou is available at (urlhttps://github.com/RuBP17/AlphaDou).
2 The Game of Doudizhu
@@ 2.1 Bidding
2.2 Cardplay
During the Cardplay phase, players take turns playing cards. Each game consists of multiple rounds, starting with a player playing a valid card combination (e.g., solo, pair). The first round is initiated by the Landlord. Subsequent players must either pass or defeat the previous hand by playing a higher-ranked combination (an action has a rank, refer to Appendix A). The round continues until two consecutive players pass. Then, the player who played the last hand starts the next round. The objective is to clear all cards from one’s hand to win. Each "Bomb" and "Rocket" can double the game’s stakes. If the Landlord wins, they receive double the rewards, whereas if the Peasant team wins, each Peasant player receives single rewards. Rewards are influenced by the bid score and the presence of Bombs or Rockets. Bombs surpass any action. The only way to defeat a Bomb is with a higher-ranked Bomb or a Rocket. The Rocket is the highest action in the game and can beat any Bomb or action. When a Bomb or Rocket is played, the points at stake double. For instance, if the winning bid is 3 points at the start, it becomes 6 points if a Bomb is played and 12 points if another Bomb is played. With two Bombs played, the Landlord stands to win/lose 24 points, and each Peasant stands to win/lose 12 points. The game concludes when a player clears all their cards. To encourage more aggressive play, Doudizhu includes a "Spring" reward: if throughout the game, the Peasant team makes no plays other than passing, or the Landlord only passes once, it is termed as Spring or Anti-Spring, respectively, doubling the reward (equivalent to playing an additional Bomb).
@@ 3 AlphaDou
3.1 Card Representation and Neural Architecture
@@
3.2 Deep Monte-Carlo
Monte Carlo (MC) methods are a class of methods that estimate strategies and value functions by modeling sample paths. Monte Carlo methods are very effective in episodic tasks to estimate the value function by taking every-visit MC approach (Sutton, 1998).
@@ 3.2 Deep Monte-Carlo
3.3 DMC with Probability and Value Factorization
Considering the significant gap between the distribution of winning scores and losing scores, we perform Value Factorization on the Q-value (Wang et al., 2023). Given that there are no ties in Doudizhu and the outcomes of the game (win or lose) are mutually exclusive, we have:
@@ 3.3 DMC with Probability and Value Factorization
We use the epsilon-greedy method to introduce exploration into the strategy $\pi(s)$ used to generate data. In appendix B, we utilized the win rate model douzero-wp and the expected value model douzero-adp provided by DouZero to verify that our proposed strategy generation method outperforms the "choosing the move with the highest expectation" strategy.
4 EXPERIMENTS
In this chapter, we compare the performance of the AlphaDou card-playing model (CardModel) with Douzero and Douzero Resnet. Douzero Resnet is the state-of-the-art Doudizhu AI based on the Douzero algorithm, replacing the LSTM neural network in Douzero with ResNet, significantly improving performance compared to Douzero. The weights and code for Douzero Resnet are open-sourced at https://github.com/Vincentzyx/Douzero_Resnet. We also compare the AlphaDou bid model (Bid Model) with a supervised learning bid model (Douzero Resnet Bid). The Douzero Resnet Bid we used is derived from Douzero Resnet: fixing the landlord player’s hand, randomly distributing 1000 sets of farmer hands and landlord hands, loading the Douzero Resnet model for games, obtaining a mean score from the results of 1000 games, using the hand as input, and the mean score as the label for supervised learning. When applying the model, a threshold is set for bidding: a model output greater than -0.1 bids 1 point, greater than 0 bids 2 points, and greater than 0.1 bids 3 points. The Douzero demonstration website https://www.douzero.org/bid also has a bidding model, but its bidding method is not based on a 3-point system, and it does not provide models for tests. Our AI system is trained on a server with 4 Intel(R) Xeon(R) Gold 6330 CPUs @ 2.10GHz and a GeForce RTX 4090 GPU in the Ubuntu 20.04 operating system.
4.1 Compare Bid Model to Douzero Resnet Bid
@@
Metrics. Following (Jiang et al., 2019), given an algorithm A and an opponent B, we use two metrics to compare the performance of A and B:
@@
In the Appendix C, we provide a detailed analysis of the performance of the Bid Model in real gameplay scenarios, illustrating the strategic improvements it offers compared to Douzero Resnet Bid. The Bid Model demonstrates the ability to adjust its strategy based on the opponent’s actions and may also adopt a more aggressive bidding approach to force the opponent into retreat (bluffing).
4.2 Compare Card Model to Benchmarks with Random Bidding Phase
To evaluate the performance of the Card Model, we followed the approach of (Jiang et al., 2019) and Douzero (Zha et al., 2021), initiating a competition between the Landlord and the Peasants. We reduce variance by playing each deck twice. Specifically, for two competing algorithms A and B, they will first play with A as the Landlord and B as the Peasants for a given deck. Then, they swap roles, with A as the Peasants and B as the Landlord, and play the same deck again. A total of 4,000 games were conducted. Considering that the Bid result is random, we set the initial score of the game to 2 points for the Landlord’s win and 1 point for the Peasants’ win, with each bomb doubling the final score (we define this scoring method as ADP1). The Card Model needs to decide the playing strategy based on the bidding process, and the random bidding process will lead to a decline in model performance because the random testing deck distribution deviates from the training process.
@@ A B Card Model Douzero Resnet Douzero
Card Model - - 0.522 0.103 0.597 0.434
Douzero Resnet -0.478 -0.103 - - 0.570 0.269
Douzero 0.403 -0.434 0.423 -0.269 - -
@@
4.3 Compare Card Model to benchmarks with Bidding Phase
@@
Card Model - - - 0.544 0.315 0.738 0.617 0.576 1.585
Douzero Resnet 0.456 -0.315 -0.738 - - - 0.581 0.314 0.937
Douzero 0.383 -0.576 -1.585 0.419 -0.314 -0.937 - - -
@@
Card Model 0.514 -0.048 0.785 4.645
Douzero Resnet 0.490 -0.209 0.771 4.296
@@
In the Appendix D, we provide a detailed analysis of the performance of the Card Model in real gameplay scenarios. Compared to Douzero and Douzero Resnet, the Card Model is more adept at accurately assessing the current state, making it better at predicting the opponent’s hand and discerning their intentions.
5 Conclusion